# On completeness of Foliated structures, and Null Killing fields

### v1

@@ 4.3. The foliated problem
Here we construct an example (Example 4.5) where Problem 1.8 fails. The key fact that makes the problem to fail is the existence of compact manifolds with an incomplete foliated flat Lorentzian structure (Example 3.3).
Example 4.4 (Compact Lorentz manifold with incomplete flat Lorentzian leaves).
@@
This is in fact a (non-flat) plane wave metric on ${\mathsf{Sol}}$ (normal-(((see [4, Theorem 1.4])normal-))). The left invariant distribution induced by ${\sf{span}}(A,X)$ is integrable, and a simple computation shows that it is flat, so that the induced metric is a left invariant Lorentzian flat metric on $\sf{Aff}(\mathbb{R})$, which is then incomplete. In fact, the above metric on ${\mathsf{Sol}}$ is given explicitly in coordinates in [4, Theorem 1.4], and the incomplete flat foliation there corresponds to the levels of the $x$-coordinate. Now, consider the left quotient of ${\mathsf{Sol}}$ by some lattice $\Gamma$. The left action of $\Gamma$ preserves this foliation, which is then well defined in the compact quotient $\Gamma\backslash{\mathsf{Sol}}$. Its leaves are incomplete, and develop into the half Minkowski plane. This gives an example of an incomplete foliated $(G,X)$-structure, where the geometry here is the flat Lorentzian structure. It is known that this structure is complete when the manifold is compact (Carrière theorem). This example shows in particular that a foliated Carrière theorem is not true.
Example 4.5 (Incomplete foliated $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-structure).
Consider the product ${\mathsf{Sol}}\times\mathbb{R}$, with the basis $(A,X,Y,V)$ of the Lie algebra, where $V$ is the basis for the $\mathbb{R}$-factor. Define a left invariant Lorentzian metric on it such that: ${\mathsf{Sol}}$ is endowed with the metric given in the previous example, $\mathbb{R}$ is Riemannian, and the two spaces are orthogonal. Namely, in the basis $(A,X,Y,V)$
@@
This is a subgeometry (stiffened) of the $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-geometry. The previous example shows that the leaves are incomplete. This gives an example of an incomplete foliated $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-structure. This example generalizes to a ${\sf{L}}_{\sf{u}}(n-1,1)$-foliated structure.
Remark 4.6 (Incomplete foliated $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$-structure).
The example above shows that Problem 1.8 also fails for the dual $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$-geometry. Indeed, up to permute the basis, the structure group becomes
@@
which is also a subgeometry of $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$.
@@ 5.1. Case of a totally geodesic lightlike foliation of a Lorentzian manifold
Let $(M,\mathfrak{F})$ be a compact Lorentzian manifold with a totally geodesic lightlike foliation $\mathfrak{F}$. Each leaf has a degenerate Riemannian metric with $1$-dimensional kernel: we denote by $\mathcal{N}$ the foliation tangent to the kernel.
@@
Let $(M,\mathfrak{F})$ be a compact Lorentzian manifold with a codimension $1$ totally geodesic lightlike foliation $\mathfrak{F}$. Then the leaves of $\mathfrak{F}$ are complete, with respect to the induced connection, if and only if the (null) geodesics tangent to $\mathcal{N}$ are complete.
Geodesic equation
Near any point $p\in M$, there exists local coordinates $(v,u,x=(x_{1},...,x_{n}))$ where the metric has the following form (see [5, Proposition 2.2]):
$g=2dudv+H(v,u,x)du^{2}+\sum_{i=1}^{n}W_{i}(v,u,x)dudx^{i}+\sum_{i,j}h_{ij}(u,x)dx^{i}dx^{j}.$
Let $V:=\partial_{v}$. In those coordinates, we have $\nabla_{V}V=0$, so the flow of $V$ parameterizes the null geodesics in $\mathcal{N}$. The leaves of $\mathfrak{F}$ are given by the levels of $u$.
For a geodesic tangent to $\mathfrak{F}$, the $u$-coordinate is constant along the geodesic, so that $\dot{u}=0$. For such a geodesic, the geodesic equation simplifies into the following
@@
We see that the coefficients in the geodesic equations are time independent.
Let $E$ be a smooth $n$-dimensional distribution on $M$, such that for any $p\in M$, $E_{p}$ is a non-degenerate Riemannian subspace of $T_{p}M$ tangent to $\mathfrak{F}$. For every $p\in M$, there exists an open neighborhood $O_{p}$ of 00 in $E_{p}$ such that the exponential map $\exp_{p}:E_{p}\subset T_{p}M\to M$ restricted to $E_{p}$ is a diffeomorphism from $O_{p}$ onto its image in $M$. Set $S_{p}:=\exp_{p}O_{p}$; it is an $n$-submanifold through $p$. For every $p\in M$, consider a normal coordinate system $(x^{1},...,x^{n})$ on $S_{p}$. If $\rho=\sum(x^{i})^{2}$, then for $\delta>0$ sufficiently small, $B_{p}(\delta)=\{q\in S_{p},\rho(q)<\delta\}$ is a normal neighborhood of $p$ in $S_{p}$ diffeomorphic to an open ball $B_{n}(0,\delta)\subset\mathbb{R}^{n}$.
Since $M$ is compact, a continuity argument shows that there exists $r>0$ such that for any $p\in M$, we can take $B_{p}$ above to be the open ball of radius $r$ and center $p$ in $S_{p}$. Now, by compactness of $M$ and the fact that the null geodesics in the kernel of $\mathfrak{F}$ are complete, there exists a positive constant, which we denote again by $r$, such that every point $p\in M$ admits a coordinate chart $(u,v,x^{1},...,x^{n}):U_{p}\subset M\to I\times J\times B_{n}(0,r)$ defined on an open neighborhood $U_{p}$ of $p$, such that $|I|=|J|=r$. Henceforth, to each point $p\in M$ we associate a coordinate chart as before. We can assume without loss of generality that $r=1$. Denote by $h_{u}$ the Riemannian metric induced on the slice $\{u\}\times\{v\}\times B_{n}(0,r)$. It does not depend on $v$.
@@
Let $\gamma$ be a geodesic tangent to $\mathfrak{F}$, such that $\gamma(0)=p\in M$. Let $(u,v,x)$ be a system of coordinates around $p$ as above. Then the $h_{u}$-norm $||\dot{x}(t)||_{u}$ is independent of the coordinate system, since it coincides with $g(\dot{\gamma}(t),\dot{\gamma}(t))$. Moreover, the second order differential equation on $x(t)$ is exactly the geodesic equation for the Riemannian metric $h_{u}$, where $u:=u(p)$.
@@
Let $\gamma$ be a geodesic tangent to $\mathfrak{F}$, such that $\gamma(0)=p\in M$. Let $(u,v,x)$ be a system of coordinates around $p$ as above. If $x(t)$ is defined on $[0,\alpha]$, then $\gamma(t)$ is defined on $[0,\alpha]$.
@@
From the equation on $v(t)$, one can see that $v(t)$ is defined as soon as $x(t)$ is defined. However, given some coordinate chart $I\times J\times B_{n}(0,1)$, the $v(t)$-coordinate may well leave the interval $I$ before $x(t)$ leaves the ball $B_{n}(0,1)$. Suppose this is the case. One can slide over the kernel of $\mathfrak{F}$ and consider a new coordinate chart containing the geodesic leaving the first one. Remember that the flow of any vector field tangent to the kernel is isometric with respect to the Riemannian metric induced in the $\mathfrak{F}$-leaf containing $\gamma$. It follows that in this new chart, the equation on $x(t)$ is the same as in the first coordinate chart. One can do this until the $x$-coordinate leaves the ball $B_{n}(0,1)$. So $\gamma$ is indeed defined as soon as $x(t)$ is defined. ∎
@@
[…]
@@ 6. On completeness of compact Lorentzian manifolds with a null Killing field
Let $(M,V,g)$ be a compact Lorentzian manifold with a null Killing field $V$. Recall that for any geodesic $\gamma$, we have $g(\dot{\gamma},V)=C$ constant along $\gamma$, and $C$ is called the Clairaut’s constant. In this section, we deal with the global completeness problem, i.e. completeness of the whole manifold $M$. We have the following result from the previous section
@@ 6.1. Left invariant metrics on Lie groups
Let $G$ be a connected Lie group endowed with a left invariant pseudo-Riemannian metric, $i.e$ its Lie algebra ${\mathfrak{g}}$ is endowed with a non-degenerate inner product $q$. Given any curve $\gamma$ in $G$, we can consider the associated curve $y(t):=dL_{\gamma(t)}^{-1}(\dot{\gamma}(t))\in{\mathfrak{g}}$. By Euler-Arnold theorem, the geodesics of $G$ are in one-to-one correspondance with the integral curves of the (Euler-Arnold) quadratic-homogeneous vector field on ${\mathfrak{g}}$:
@@
Let $\mathsf{v}\in{\mathfrak{g}}$ such that $\mathrm{ad}_{\mathsf{v}}$ is skew symmetric with respect to $q$. Then the left invariant metric on $G$ is also invariant by the right multiplication by $\exp(t\mathsf{v})$. So the left invariant vector field $V$ generated by $\mathsf{v}$ is Killing, and it is everywhere isotropic. Moreover, it acts isometrically (on the right) on $\Gamma\backslash G$, where $\Gamma\subset G$ is a cocompact lattice, defining a null Killing field on the compact quotient.
@@
Indeed, one has $g(\mathrm{ad}^{*}_{y}y,v)=g(\mathrm{ad}_{y}\mathsf{v},y)=-g(\mathrm{ad}_{\mathsf{v}}y,y)=0$. ∎
@@
Let $(\mathsf{v},\mathsf{x}=(\mathsf{x_{1}},..,\mathsf{x_{n}}),\mathsf{u})$ be a basis of ${\mathfrak{g}}$ such that $(\mathsf{x},\mathsf{v})$ generate $\mathsf{v}^{\perp}$. Let $y(t)$ be an integral curve of $\dot{y}$ tangent to $\mathsf{v}^{\perp}$, and write $y(t)=(v(t),x(t),u(t))$ in the above basis. The hyperplane $\mathsf{v}^{\perp}$ is lightlike with a transverse Riemannian quadratic form. Since the geodesic vector field is tangent to the levels of the quadratic form $q$, then on $\mathsf{v}^{\perp}$, it is tangent to levels of the form $\mathbb{S}^{n}\times\mathbb{R}$, where $\mathbb{R}$ is the direction of $\mathsf{v}$ and $\dim\mathsf{v}^{\perp}=n+1$. In particular, the $x$-component is bounded, hence complete. On the other hand, one can check that $\mathrm{ad}^{*}_{v}v=0$; as a consequence, Equation (1) on $y(t)$ yields an equation on $\dot{v}$ which is linear in $v$, hence $v(t)$ is also complete. ∎

### v2

@@ 4.3. The foliated problem
Here, we construct examples where leafwise completeness fails for the aforementioned geometries.
Example 4.4 (Compact manifold with incomplete flat Lorentzian leaves).
@@
This is in fact a (non-flat) plane wave metric on ${\mathsf{Sol}}$ ((((see [3, Theorem 1.4])))). The left invariant distribution induced by ${\sf{span}}(A,X)$ is integrable, and a simple computation shows that it is flat, so that the induced metric is a left invariant Lorentzian flat metric on $\sf{Aff}(\mathbb{R})$, which is then incomplete. In fact, the above metric on ${\mathsf{Sol}}$ is given explicitly in coordinates in [3, Theorem 1.4], and the incomplete flat foliation there corresponds to the levels of the $x$-coordinate. Now, consider the left quotient of ${\mathsf{Sol}}$ by some lattice $\Gamma$. The left action of $\Gamma$ preserves this foliation, which is then well defined in the compact quotient $\Gamma\backslash{\mathsf{Sol}}$. Its leaves are incomplete, and develop into the half Minkowski plane. This gives an example of an incomplete foliated $(G,X)$-structure, where the geometry here is the flat Lorentzian structure. It is known that this structure is complete when the manifold is compact (Carrière theorem). This example shows in particular that a foliated Carrière theorem is not true.
Example 4.5 (Compact incomplete $({\sf{L}}_{\sf{u}}(p,q),\mathbb{R}^{p+q+1})$-foliated manifold).
Consider the product ${\mathsf{Sol}}\times\mathbb{R}$, with the basis $(A,X,Y,V)$ of the Lie algebra, where $V$ is the basis for the $\mathbb{R}$-factor. Define a left invariant Lorentzian metric on it such that: ${\mathsf{Sol}}$ is endowed with the metric given in the previous example, $\mathbb{R}$ is Riemannian, and the two spaces are orthogonal. So in the basis $(A,X,Y,V)$, we have
@@
This is a subgeometry (stiffened) of the $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-geometry. Now, taking a cocompact lattice of the form $\Gamma\times\mathbb{Z}\subset{\mathsf{Sol}}\times\mathbb{R}$ gives a compact, locally homogeneous, $({\sf{L}}_{\sf{u}}(1,1),\mathbb{R}^{3})$-foliated manifold. And the previous example shows that the leaves are incomplete. By taking products of the previous example, and potentially including a flat Riemannian torus, we can construct a compact ${\sf{L}}_{\sf{u}}(p,q)$-foliated manifold with incomplete leaves for any $p,q\geq 1$. Note that here, the holonomy representation of the fundamental group of any leaf preserves a Lorentzian metric, since it is a subgroup of ${\sf{O}}(1,1)\ltimes\mathbb{R}^{3}\subset{\sf{L}}_{\sf{u}}(1,1)\ltimes\mathbb{R}^{3}$ (with a non-trivial projection to the ${\sf{O}}(1,1)$-factor). Observe also that the leaves develop injectively, so the structure is Kleinian!
Remark 4.6 (Compact incomplete $({\sf{L}}_{\sf{u}}^{*}(p,q),\mathbb{R}^{p+q+1})$-foliated manifold).
The example above shows that Problem 1.8 also fails for the dual $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$-geometry. Indeed, up to permuting the basis, the structure group becomes
@@
which is also a subgeometry of $({\sf{L}}_{\sf{u}}^{*}(1,1),\mathbb{R}^{3})$. From this example, one can obtain in the same way as in Example 4.5 a compact ${\sf{L}}_{\sf{u}}^{*}(p,q)$-foliated manifold with incomplete leaves for any $p,q\geq 1$.
@@ 5.1. Case of a totally geodesic lightlike foliation of a Lorentzian manifold
Let $(M,g,\mathfrak{F})$ be a compact Lorentzian manifold with a totally geodesic lightlike foliation $\mathfrak{F}$. The metric induced on each leaf is a Riemannian degenerate metric. We denote by $\mathcal{N}$ the ($1$-dimensional) foliation tangent to the radical. The goal of this section is to prove the following result, which corresponds to the situation described in Item (A) of the introduction.
Let us mention that the foliation $\mathcal{N}$, when restricted to each leaf $F$ of $\mathfrak{F}$, is transversely Riemannian [38]. In other words, the local flow of any vector field tangent to $\mathcal{N}$ preserves the induced degenerate Riemannian metric on $F$.
@@
Let $(M,g,\mathfrak{F})$ be a compact Lorentzian manifold with a codimension $1$ totally geodesic lightlike foliation $\mathfrak{F}$. Then the leaves of $\mathfrak{F}$ are complete with respect to the induced connection if and only if the (null) geodesics tangent to $\mathcal{N}$ are complete.
5.1.1. Geodesic equation
Let $\dim M=n+2$. Near any point $p\in M$, there exist local coordinates $(u,v,x=(x^{1},\ldots,x^{n}))$ where the metric has the following form (see [5, Proposition 2.2]):
$g=2dudv+H(u,v,x)du^{2}+\sum_{i=1}^{n}W_{i}(u,v,x)dudx^{i}+\sum_{i,j}h_{ij}(u,x)dx^{i}dx^{j}.$
We will refer to them as adapted local coordinates.
Set $V:=\partial_{v}$. It is easy to check that $\nabla_{V}V=0$, so the local vector field $V$ is geodesic: its flow parametrizes the null geodesics in $\mathcal{N}$. The leaves of $\mathfrak{F}$ are given by the levels of $u$. For a geodesic tangent to $\mathfrak{F}$, the $u$-coordinate is constant along the geodesic, so we have $\dot{u}=0$. And the geodesic equation is the following
@@ 5.1.1. Geodesic equation
The fact that the coefficients in $x(t)$ do not depend on $v$ follows from the fact that the foliation $\mathcal{N}$ is transversely Riemannian.
5.1.2. Proof of Theorem 5.1
Let $E$ be a smooth $n$-dimensional distribution on $M$, such that for any $p\in M$, $E_{p}$ is a non-degenerate Riemannian subspace of $T_{p}M$ tangent to $\mathfrak{F}$. This can be defined by taking an auxiliary Riemannian metric $h$, and setting $E:=T\mathcal{N}^{\perp_{h}}\cap T\mathcal{N}^{\perp_{g}}$.
For every $p\in M$, there exists a convex neighborhood $O_{p}$ of  00  in $E_{p}$. The exponential map $\exp_{p}:E_{p}\subset T_{p}M\to M$ restricted to $E_{p}$ is a diffeomorphism from $O_{p}$ onto its image in $M$. Define the $n$-submanifold $S_{p}:=\exp_{p}O_{p}$ through $p$. For every $p\in M$, consider a normal coordinate system $(x^{1},\ldots,x^{n})$ on $S_{p}$. If $\rho=\sum(x^{i})^{2}$, then for $r_{p}>0$ sufficiently small, $B(p,r_{p})=\{q\in S_{p},\rho(q)<r_{p}\}$ is a normal neighborhood of $p$ in $S_{p}$ diffeomorphic to $B_{n}(0,r_{p})$, an open ball of $\mathbb{R}^{n}$ of center 00 and radius $r_{p}$. Then, we define adapted local coordinates in a neighborhood $U_{p}$ of $p$:
$\phi_{p}=(u,v,x=(x^{1},\ldots,x^{n})):U_{p}\to I_{p}\times J_{p}\times B_{n}(0,r_{p}),$
where $I_{p}$ and $J_{p}$ are open intervals of $\mathbb{R}$. Since $M$ is compact, a continuity argument shows that there exists $r>0$ such that for any $p\in M$, we can take $B(p,r_{p})$ to be of uniform radius $r>0$.
Observation 5.2.
$(1)$ Consider adapted local coordinates $(u,v,x)$ on $U$, with the associated coordinate vector fields $Z=\partial_{u},V=\partial_{v}$. The choice of other local coordinates on $U^{\prime}\subset U$ gives rise to coordinate vector fields $Z_{1}=\partial_{u_{1}},V_{1}=\partial_{v_{1}}$ such that, when restricted to a leaf of $\mathfrak{F}$, $V_{1}$ is a rescaling of $V$ by some constant. Indeed, since both $Z$ and $Z_{1}$ map a leaf of $\mathfrak{F}$ to a leaf of $\mathfrak{F}$, then $Z_{1}=f(u)Z+W$ for some $\mathfrak{F}$-invariant (local) function $f$ and some (local) vector field $W$ tangent to $\mathfrak{F}$. Since $V_{1}$ is determined by the equation $g(Z_{1},V_{1})=1$, the claim follows.
$(2)$ Given adapted local coordinates on $U$, for any $v$, the subset $\{0\}\times\{v\}\times B_{n}(0,r)$ is also an open ball of radius $r$ for the induced Riemannian metric on it. This follows from the fact that the foliation $\mathcal{N}$, when restricted to each leaf $F$ of $\mathfrak{F}$, is transversely Riemannian. More explicitly, the local flow of the vector field $\partial_{v}$, which is tangent to $\mathcal{N}$ in $U$, preserves the induced degenerate Riemannian metric on $F\cap U$.
We can assume without loss of generality that $r=2$.
Notation: Denote by $h_{u}$ the Riemannian metric induced on the slice $\{u\}\times\{v\}\times B_{n}(0,2)$, for some $v$ (it is independent of $v$).
@@
Let $\gamma$ be a geodesic tangent to $\mathfrak{F}$. Let $(u,v,x)$ be a system of coordinates around $\gamma(0)$, as above. Then, the $h_{u}$-norm of $\dot{x}(t)$ is independent of the coordinate system, since it coincides with $g(\dot{\gamma}(t),\dot{\gamma}(t))$. And the second order differential equation on $x(t)$ reads
$\nabla^{h_{u}}_{\dot{x}(t)}\dot{x}(t)=0.$ (1)
It is the geodesic equation associated to the Riemannian metric $h_{u}$, with $u:=u(\gamma(0))$.
@@
Assume that the null geodesics tangent to $\mathcal{N}$ are complete. Let $\gamma$ be a (maximal) geodesic tangent to $\mathfrak{F}$. If $g(\gamma^{\prime}(0),\gamma^{\prime}(0))=1$, then $\gamma(t)$ is defined on $[0,1]$.
@@
[…]
For this, let $c(t)$ be the maximal (parameterized) geodesic tangent to $\mathcal{N}$ at $p$. Let $p_{0}:=c(t_{0})$, and consider a neighborhood $U_{p_{0}}$ of $p_{0}$, on which we have adapted local coordinates $\phi_{0}=(u_{0},v_{0},x_{0}):U_{p_{0}}\to I_{0}\times J_{0}\times B_{n}(0,2)$. Since the null geodesics tangent to $\mathcal{N}$ are complete, the (smooth) slices $S_{0}(v):=\phi_{0}^{-1}(I_{0}\times\{v\}\times B_{n}(0,2))\subset U_{p_{0}}$, for $v\in J_{0}$, converge to a submanifold $B_{0}$ of $M$ which is transversal to $\mathcal{N}$. By Observation 5.2 $(2)$, the intersection of this submanifold with $F$ is an open ball of radius $2$ for the induced Riemannian metric on it. Let $p_{1}:=c(t_{1})\in B_{0}$. Denote by $Z_{0},V_{0},X^{1}_{0},\ldots,X_{0}^{n}$ the coordinate vector fields on $U_{p_{0}}$. These vector fields can be extended to coordinate vector fields $Z_{1},V_{1},X^{1}_{1},\ldots,X^{n}_{1}$ on some neighborhood $U_{p_{1}}$ of $p_{1}$, in such a way that the associated coordinates $(u_{1},v_{1},x_{1})\in I_{1}\times J_{1}\times B_{n}(0,2)$ are adapted coordinates satisfying $u_{1}=u_{0},v_{1}=v_{0},x_{1}=x_{0}$ on $U_{p_{0}}\cap U_{p_{1}}$. We say that these coordinates are compatible with those on $U_{p_{0}}$.
One can define this way a sequence of compatible local charts $U_{i}:=U_{p_{i}}$, where $p_{i}:=c(t_{i})$ is an increasing sequence of points on $c(t)$ such that $p_{0}=p$, $U_{i}\cap U_{j}\neq\emptyset$ if and only if $j=i+1$. Denote by $\phi_{i}=(u_{i},v_{i},x_{i})\in I_{i}\times J_{i}\times B_{n}(0,1)$ the local adapted coordinates on $U_{i}$. We claim that such a sequence of open sets can be obtained with $t_{i}\to\infty$. To see this, define the slices $S_{i}(v):=\phi_{i}^{-1}(I_{i}\times\{v_{i}\}\times B_{n}(0,2))\subset U_{i}$, and denote by $S(p_{i})$ the slice containing $p_{i}$. Suppose that the above sequence does not satisfy $t_{i}\to\infty$. Since $c(t)$ is complete, the points $p_{i}$ converge to a point $p\in c(t)$, and the slices $S(p_{i})$ (which are leaves of the foliation defined by the levels of $v_{i}$ in $U_{i}$) converge to a submanifold $S(p)$ through $p$. The intersection of this submanifold with $F$ is an open ball of radius $2$ for the induced Riemannian metric on it. Define adapted coordinates $(u,v,x)$ in a neighborhood $U$ of $p$, such that the slice through $p$ in $U$ given by the level of $v$ at $p$ coincides with $S(p)$. We call this ‘a good overlapping property’. Set $V=\partial_{v}$. There exists some $U_{n}$ such that $U_{n}\cap U\neq 0$. By Observation 5.2 $(1)$, the restriction of $V$ to a leaf of $\mathfrak{F}$ is a rescaling of the restriction of $V_{n}$ to that leaf by some constant. Due to this and to the good overlapping property, the slices defined by the levels of $v$ in $U\cap U_{n}$ coincide with those defined by the levels of $v_{n}$. Thus, we can define new adapted coordinates on $U$ that are compatible with those on $U_{n}$, thereby extending the previous sequence beyond $p$.
By construction, the geodesic equations associated to two overlapping open sets $U_{i}$ and $U_{i+1}$ coincide on $U_{i}\cap U_{i+1}$, giving rise to a single geodesic equation:
$\ddot{x}^{k}(t)+\Gamma_{ij}^{k}(0,x)\,\dot{x}^{i}\dot{x}^{j}=0$
$\ddot{v}(t)+\Gamma_{ij}^{u}(0,v,x)\,\dot{x}^{i}\dot{x}^{j}=0$
$u=0$
where the coefficients are, this time, defined on $\{0\}\times\mathbb{R}\times B$, with the closure of $B_{n}(0,1)$ contained in $B$. The fact that $v$ is defined on $\mathbb{R}$ follows from the fact that the geodesics tangent to $\mathcal{N}$ are complete.
Thus, to the geodesic $\gamma_{|\bigcup_{i}U_{i}}$ corresponds a maximal geodesic $\widetilde{\gamma}$ in $\{0\}\times\mathbb{R}\times B$, defined on the same interval as $\gamma_{|\bigcup_{i}U_{i}}$, and satisfying the geodesic equation above. For this geodesic $\widetilde{\gamma}$, it is clear that $x(t)$ is defined on $[0,1]$, and $v(t)$ is defined on the same interval as $x(t)$. Consequently, $\widetilde{\gamma}$, and therefore $\gamma$, is defined on $[0,1]$. ∎
@@
This follows from the lemma above, and the fact that the $g$-norm of $\gamma^{\prime}(t)$ is preserved on $\gamma$. ∎
@@ 6. On completeness of compact Lorentzian manifolds with a null Killing field
Let $(M,g,V)$ be a compact Lorentzian manifold with a null Killing field $V$. Recall that for any geodesic $\gamma$, we have $g(\dot{\gamma},V)=C$ constant along $\gamma$, and $C$ is called the Clairaut’s constant. In this section, we deal with the global completeness problem, i.e. completeness of the whole manifold $M$. We have the following result from the previous section
@@ 6.1. Left invariant metrics on Lie groups
Let $G$ be a connected Lie group endowed with a left invariant pseudo-Riemannian metric, $i.e$ its Lie algebra ${\mathfrak{g}}$ is endowed with a non-degenerate inner product $q$. Given any curve $\gamma$ in $G$, we can consider the associated curve $y(t):=dL_{\gamma(t)}^{-1}(\dot{\gamma}(t))\in{\mathfrak{g}}$. By Euler-Arnold theorem, the geodesics of $G$ are in one-to-one correspondence with the integral curves of the (Euler-Arnold) quadratic-homogeneous vector field on ${\mathfrak{g}}$:
@@
Let $\mathsf{v}\in{\mathfrak{g}}$ such that $\mathrm{ad}_{\mathsf{v}}$ is skew symmetric with respect to $q$. Then the left invariant metric on $G$ is also invariant by the right multiplication by $\exp(t\mathsf{v})$. So the left invariant vector field $V$ generated by $\mathsf{v}$ is Killing, and it is everywhere null. Moreover, it acts isometrically (on the right) on $\Gamma\backslash G$, where $\Gamma\subset G$ is a cocompact lattice, defining a null Killing field on the compact quotient.
@@
Indeed, one has $g(\mathrm{ad}^{*}_{y}y,\mathsf{v})=g(\mathrm{ad}_{y}\mathsf{v},y)=-g(\mathrm{ad}_{\mathsf{v}}y,y)=0$. ∎
@@
Let $(\mathsf{v},\mathsf{x}=(\mathsf{x_{1}},\ldots,\mathsf{x_{n}}),\mathsf{u})$ be a basis of ${\mathfrak{g}}$ such that $(\mathsf{v},\mathsf{x})$ generates $\mathsf{v}^{\perp}$. Let $y(t)$ be an integral curve of $\dot{y}$ tangent to $\mathsf{v}^{\perp}$, and write $y(t)=(v(t),x(t),u(t))$ in the above basis. The hyperplane $\mathsf{v}^{\perp}$ is lightlike with a transverse Riemannian quadratic form. Since the geodesic vector field is tangent to the levels of the quadratic form $q$, then on $\mathsf{v}^{\perp}$, it is tangent to levels of the form $\mathbb{S}^{n}\times\mathbb{R}$, where $\mathbb{R}$ is the direction of $\mathsf{v}$ and $\dim\mathsf{v}^{\perp}=n+1$. In particular, the $x$-component is bounded, hence complete. On the other hand, one can check that $\mathrm{ad}^{*}_{\mathsf{v}}\mathsf{v}=0$; as a consequence, Equation (2) on $y(t)$ yields an equation on $\dot{v}$ which is linear in $v$, hence $v(t)$ is also complete. ∎